the members of a gardening club decided to spend $36 on buying garden tools and they found that a rake and a travel together cost $7. If they spent all the money on rakes alone, they could buy 3 more tools than they could if they spent all the money on trowels.
a) By taking the cost of atrowel as $q, write down expression in terms of q for i) the cost of a rake
ii) the number of rakes which could be bought for $36.
b) write down an equation which q must satisfy, and show that it reduces to q^2+17q-84=0.

Answer by Theo(13342)


let R = cost of a rake.
let Q = cost of a trowel.

you are given that the cost of a rake plus the cost of a trowel = 7.

that gets you R + Q = 7

solve for the cost of a rake to get R = 7-Q

let x equal the number of rakes that can be bought for 36 dollars.

that gets you x * R = 36

since R = 7-Q, that gets you x * (7-Q) = 36

solve for x to get x = 36 / (7-Q)

x is the number of rakes that you can buy for 36 dollars when the cost of a rake is equal to 7-Q.

if you buy all rakes or all trowels, you can buy 3 more rakes than trowels.

this means you can buy either x rakes or (x-3) trowels for 36 dollars.

since the cost of a rake is (7-Q) and the cost of a trowel is Q, you get x * (7-Q) = 36 and you get (x-3) * Q = 36

since they're both equal to 36, then they're both equal to each other and you get x * (7-Q) = (x-3) * Q.

since x * (7-Q) = 36, then x * (7-Q) = (x-3) * Q becomes 36 = (x-3) * Q

since x = 36 / (7-Q), then 36 = (x-3) * Q becomes 36 = ((36 / (7-Q)) - 3) * Q

multiply both sides of this equation by (7-Q) to get 36 * (7-Q) = (36 - 3 * (7-Q) * Q

simplify this to get 252 - 36Q = (36 - 21 + 3Q) * Q

simplify this further to get 252 - 36Q = 36Q - 21Q + 3Q^2

re-order the terms in descending order of degree to get -36Q + 252 = 3Q^2 + 36Q - 21Q

combine like terms to get -36Q + 252 = 3Q^2 + 15Q

add 36Q to both sides of this equation to get 252 = 3Q^2 + 15Q + 36Q

combine like terms to get 252 = 3Q^2 + 51Q

subtract 252 from both sides of this equation to get 0 = 3Q^2 + 51Q - 252

factor a 3 out of the expression on the right side of the equation to get 0 = Q^2 + 17Q - 84

this is the same as Q^2 + 17Q - 84 = 0

that is the equation you are looking for.

if you factor this equation, you will get (Q + 21) * (Q - 4) = 0

solve for Q to get Q = 4 or Q = -21.

since Q can't be negative, then Q = 4.

you were given that R + Q = 7

when Q = 4, you get R = 3

the price of a rake is 3 and the price of a trowel is 4.

36/3 gets you 12 rakes.

36/4 gets you 9 trowels.

12 is 3 greater than 9 which means the number of rakes you can buy with 36 dollars is 3 more than the number of trowels you can buy with 36 dollars.

all the requirements of the problem are satisfied.

the answers to your problem are:

ai) the cost of a rake is equal to 7-Q.

aii) the number of rakes which could be bought for $36 is equal to 36/(7-Q).

b) the equation is x * (7-Q) = (x-3) * Q, where x is the number of rakes and (x-3) is the number of trowels.

the equation was reduced to Q^2 + 17Q - 84 = 0 as shown above.